OpenStudy (anonymous):

A car worth $3600 is taxed $201. At that rate, what is the tax on a car worth $2400?

OpenStudy (gw2011):

Set up a proportion: 201/3600=x/2400 Cross multiply: 3600x=(201)(2400) 3600x=482,400 x=$134 tax

OpenStudy (anonymous):

Thanks man

OpenStudy (gw2011):

You're welcome

OpenStudy (anonymous):

2400x201/3600=134 tax.